OpenStudy (anonymous):

Use the property of exponents to simplify each expression \[(16x^4)_{2}^{3}\]

OpenStudy (anonymous):

supposed to be 3/2

OpenStudy (anonymous):

i am going to make a guess that this is \[(16x^4)^{\frac{3}{2}}\] is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Exponents in that format can be multiplied together

OpenStudy (anonymous):

So youve got 16^12/2

OpenStudy (anonymous):

\[16x ^{12/2}\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

or 16x^6

OpenStudy (anonymous):

ok then we write \[16^{\frac{3}{2}}(x^4)^{\frac{3}{2}}\] then \[16^{\frac{3}{2}}=\sqrt{16}^3=4^3=64\] and and \[(x^4)^{\frac{3}{2}}=x^{4\times \frac{3}{2}}=x^6\]

OpenStudy (anonymous):

hi satellite73

OpenStudy (anonymous):

hello for a final answer of \[64x^6\]

OpenStudy (anonymous):

btw it is not \(16x^6\) because you have to raise 16 to the 3/2 power as well, since it is inside parentheses

OpenStudy (anonymous):

sorry :P

OpenStudy (anonymous):

so you just combine the two parts together for the final answer

OpenStudy (anonymous):

yes, final answer is \(64x^6\)

OpenStudy (anonymous):

ty satellite :)